Two cars located 20 miles away from each other start (at the same time) driving toward each other. One car is driving at 35 mph, the second car is driving at 45 mph. How long will it take for these two cars to meet?
Answer by TimothyLamb(4379)

relative speed of the two cars = 35 + 45 = 80
s = d/t
t = d/s
t = 20/80
t = 1/4 hour
t = 15 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php